# OmniGen: Unified Image Generation

### v1

@@ Abstract
The emergence of Large Language Models (LLMs) has unified language generation tasks and revolutionized human-machine interaction. However, in the realm of image generation, a unified model capable of handling various tasks within a single framework remains largely unexplored. In this work, we introduce OmniGen, a new diffusion model for unified image generation. Unlike popular diffusion models (e.g., Stable Diffusion), OmniGen no longer requires additional modules such as ControlNet or IP-Adapter to process diverse control conditions. OmniGen is characterized by the following features: 1) Unification: OmniGen not only demonstrates text-to-image generation capabilities but also inherently supports various downstream tasks, such as image editing, subject-driven generation, and visual-conditional generation. Additionally, OmniGen can handle classic computer vision tasks by transforming them into image generation tasks, such as edge detection and human pose recognition. 2) Simplicity: The architecture of OmniGen is highly simplified, eliminating the need for additional text encoders. Moreover, it is more user-friendly compared to existing diffusion models, enabling complex tasks to be accomplished through instructions without the need and cost for extra preprocessing steps (e.g., human pose estimation), thereby significantly simplifying the workflow of image generation. 3) Knowledge Transfer: Benefit from learning in a unified format, OmniGen effectively transfers knowledge across different tasks, manages unseen tasks and domains, and exhibits novel capabilities. We also explore the model’s reasoning capabilities and potential applications of chain-of-thought mechanism. This work represents the first attempt at a general-purpose image generation model, and there remain several unresolved issues. We will open-source the related resources at https://github.com/VectorSpaceLab/OmniGen to foster advancements in this field.
Figure 1: OmniGen demonstrates the capability to perform various image generation tasks within a single framework. Additionally, it possesses reasoning abilities and in-context learning capabilities.
@@ 1 Introduction
The pursuit of Artificial General Intelligence (AGI) has intensified the demand for generative foundation models capable of handling a wide variety of tasks within a single framework. In the field of Natural Language Processing (NLP), Large Language Models (LLMs) have become exemplary in achieving this goal, demonstrating remarkable versatility across numerous language tasks such as question answering, text summarization, and code generation.
However, the field of visual generation has yet to reveal a counterpart that mirrors the universality of LLMs. Current image generation models have demonstrated proficiency in specialized tasks. For instance, in the text-to-image generation filed, state-of-the-art models such as the Stable Diffusion series [56; 52; 13], DALL-E [55], and Imagen [26] have made significant strides. Meanwhile, many efforts have been proposed to extend and optimize the capabilities of diffusion models for specific tasks. Models like ControlNet [73] and T2i-Adapter [45] design an additional network plugged into the text-to-image diffusion model to support visual conditions. InstructPix2Pix [4] is trained on a comprehensive dataset tailored for image editing tasks. Despite their strengths, those models are limited by their task-specific nature and do not exhibit the comprehensive perceptual understanding and generative capabilities required for a universal model in visual generation.
Is it possible to address various image generation tasks, such as text-to-image, image editing, controllable generation, and image restoration, within a single diffusion framework, akin to how GPT handles language tasks? If a universal model is available, the need for training additional modules (e.g., ControlNet, IP-Adapter, T2I-Adapter) in practical applications can be eliminated. Motivated by this potential, we explore a unified framework for image generation, named OmniGen.
@@
Unlike popular diffusion models, OmniGen features a very concise structure, comprising only two main components: a VAE and a transformer model, without any additional encoders. OmniGen supports arbitrarily interleaved text and image inputs as conditions to guide image generation, rather than text-only or image-only conditions. To train a robust unified model, we construct the first large-scale unified image generation dataset X2I, which unifies various tasks into one format. Additionally, we incorporate several classic computer vision tasks such as human pose estimation, edge detection, and image deblurring, thereby extending the model’s capability boundaries and enhancing its proficiency in complex image generation tasks. We evaluate our model on multiple benchmarks, demonstrating its competitive text-to-image generation capabilities compared to existing models. Furthermore, our model inherently supports various image generation tasks, such as image editing, visual conditional generation, and subject-driven generation, which are beyond the reach of current diffusion models. Remarkably, the design of OmniGen allows for robust transfer learning across different scenarios, facilitating the handling of previously unseen tasks and domains, as well as giving birth to emerging abilities. Our contributions are summarized below:
@@
  • We introduce OmniGen, a unified model for image generation that excels in multiple domains. OmniGen demonstrates competitive text-to-image generation capabilities and inherently supports a variety of downstream tasks such as controllable image generation and subject-driven generation. Furthermore, it is capable of performing classic computer vision tasks. To the best of our knowledge, OmniGen is the first image generation model to achieve such a comprehensive level of functionality.
@@
  • We construct a comprehensive image generation dataset named X2I, which stands for "anything to image". This dataset includes a wide range of image generation tasks, all standardized into a unified format.
@@
  • By unified training on the multi-task dataset, OmniGen can apply learned knowledge to tackle unseen tasks and domains, as well as exhibit new capabilities. Additionally, OmniGen shows a degree of reasoning capability.
The remainder of this paper is organized as follows: Section 2 details the model architecture, while Section 3 describes the dataset construction. Section 4 presents the model’s performance on various image generation tasks. In Section 5, we analyze the model’s emerging capabilities and reasoning abilities, and also explore the potential applications of CoT in image generation. Section 6 discusses the current limitations of the model. Section 7 reviews related work.
@@ 2.1 Model Design
Principles. Current diffusion models are typically limited to common text-to-image tasks and can not perform a broader range of downstream image-generation tasks. To achieve real-world applications, users often need to design and integrate additional network structures to extend the capabilities of diffusion models, making the models highly cumbersome. Even worse, these additional parameter networks are usually task-specific and can not be reused for other tasks, unless more networks are designed and trained for different functions. To circumvent these issues, the design principles of OmniGen are as follows: 1). Universality: accepting any form of image and text inputs for various tasks; 2). Conciseness, avoiding overly complex structural designs and numerous additional components.
@@
Network Architecture. As illustrated in Figure 2, the OmniGen framework adopts an architecture comprised of a Variational Autoencoder (VAE) [28] and a pre-trained large transformer model. Specifically, VAE extracts continuous visual features from images, while the transformer model generates images based on input conditions. In this paper, we use the VAE from SDXL [52] and freeze it during training. We use Phi-3 [1] to initialize the transformer model, inheriting its excellent text processing capabilities. Unlike state-of-the-art diffusion models that require additional encoders to pre-process conditional information (such as clip text encoder and image encoder), OmniGen inherently encodes conditional information by itself, significantly simplifying the pipeline. Furthermore, OmniGen jointly models text and images within a single model, rather than independently modeling different input conditions with separate encoders as in existing works [67; 68; 70; 63; 9] which lacks interaction between different modality conditions.
@@
Input Format. The input to the model can be multimodal interleaved text and images in free form. We utilize the tokenizer of Phi-3 to process text without any modifications. For images, we firstly employ a VAE with a simple linear layer to extract latent representations. Then, they are flattened into a sequence of visual tokens by linearly embedding each patch in the latent space. Following [50], we apply standard frequency-based positional embeddings to input visual tokens, and use the same method as SD3 [13] to process images with varying aspect ratios. In addition, we encapsulate each image sequence with two special tokens: “<img>” and “</img>” before inserting it into the text tokens sequence. We also add the timestep embedding [50] at the end of the input sequence.
@@
Attention Mechanism. Different from text, which can be decomposed into discrete tokens to model, we argue that images should be modeled as a whole. Therefore, we modify the common causal attention mechanism in LLM, integrating it with the bidirectional attention as illustrated in Figure 2. Specifically, we apply causal attention to each element in the sequence, but apply bidirectional attention within each image sequence. This allows each patch to pay attention to other patches within the same image, while ensuring that each image can only attend to other images or text sequences that have appeared previously. .
@@
Inference. During inference, we randomly sample a Gaussian noise and then apply the flow matching method to predict the target velocity, iterating multiple steps to obtain the final latent representation. Finally, we use a VAE to decode the latent representation into the predicted image. The default inference step is set to 50. Thanks to the attention mechanism, OmniGen can accelerate inference like LLMs by using kv-cache: storing previous and current key and value states of the input conditions on the GPU to compute attention without redundant computations.
@@ 2.2 Training Strategy
Train objective. In this work, we use rectified flow [41] to optimize the parameters of model. Different from DDPM [25], flow matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the step $t$, $\mathbf{x}_{t}$ is defined as
$\mathbf{x}_{t}=t\mathbf{x}+(1-t)\boldsymbol{\epsilon},$
where $\mathbf{x}$ is the original data, and $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised data $\mathbf{x}_{t}$, timestep $t$, and condition information $c$. Specifically, the objective is to minimize the mean squared error loss:
@@
For image editing tasks, the objective is to modify specific regions of the input image while keeping other areas unchanged. Therefore, the difference between the input image and the target image is often small, which allows the model to learn an unexpected shortcut: simply copying the input image as the output to make the related training loss very low. To mitigate this phenomenon, we amplify the loss in the regions of the image where changes occur. More specifically, we calculate the loss weights for each region based on these latent representations of input image $\mathbf{x^{\prime}}$ and target image $\mathbf{x}$:
@@
Consequently, regions with alterations are assigned higher weights than those without changes, guiding the model to focus on the areas to be modified.
Training Pipeline. Following previous work [13; 18; 6], we gradually increase the image resolution during the training process. Low resolution is data-efficient, while high resolution can enhance the aesthetic quality of the generated images. Detailed information for each training stage is presented in Table 1. We adopt the AdamW [42] with $\beta=(0.9,0.999)$ as the optimizer. All experiments are conducted on 104 A800 GPUs.
Stage Image Resolution Training Steps (K) Batch Size Learning Rate
@@ 5 Multiple 80 104 2e-5
Table 1: Detailed information about every OmniGen training stage.
@@ 3 X2I Dataset
To achieve robust multi-task processing capabilities, it is essential to train models on large-scale and diverse datasets. However, in the field of image generation, a readily available large-scale and diverse dataset has yet to emerge. In this work, we have constructed a large-scale unified image generation dataset for the first time, which we refer to as the X2I dataset, meaning "anything to image". We have converted these data into a unified format, and Figure 3 presents some examples from the X2I dataset. The entire dataset comprises approximately 0.1 billion images. We will provide a detailed description of the composition of this dataset in the following sections.
Figure 3: Examples of our training data for the OmniGen model. We standardized the input of all tasks into an arbitrarily interleaved image-text sequence format, used as the model’s prompt. The placeholder |image_i| represents the position of the $i$-th image within the prompt.
@@ 3.1 Text to Image
The input for this subset of data is plain text. We have obtained multiple open-source datasets from various sources: Recap-DataComp [35](a subset of 56M images), SAM-LLaVA [6], ShareGPT4V [7], LAION-Aesthetic [58](a subset of 4M images), ALLaVA-4V [5], DOCCI [47], DenseFusion [36] and JourneyDB [60]. While these datasets are large in quantity, their image quality is not always high enough. In the early stages of training, we use them to learn a broad range of image-text matching relationships and diverse knowledge. After stage 3, we utilize our internal collection of 16 million high-quality images to enhance the aesthetic quality of the generated images. A lot of studies [13; 6] have demonstrated that synthetic detailed captions can greatly improve text-to-image models trained at scale. Therefore, we use the InternVL2 [11] to create synthetic annotations for internal data and LAION-Aesthetic (the other datasets come with detailed text descriptions and do not require further annotation).
@@ 3.2 Multi-modal to Image
Different from most existing diffusion models, our model can accept more general and flexible multimodal instruction as conditions to guide the generation of images.
@@ 3.2.1 Common Mixed-modal Prompts
The input of this portion of data is arbitrarily interleaved text and images. We collect the data from multiple tasks and sources: image editing (SEED-Data-Edit [19], MagicBrush [72], and InstructPix2Pix [4]), human motion (Something-Something [23]), virtual try-on (HR-VITON [31] and FashionTryon [75]), and style transfer (stylebooth [24]). We standardize all tasks into the input-output pair format as shown in Figure 3-(b).
The issue of utilizing additional visual conditions for finer-grained spatial control has garnered widespread attention[73; 33]. We employ the MultiGen [53] dataset to learn this function, and select six representative visual conditions: Canny, HED, Depth, Skeleton, Bounding Box, and segmentation. These types of tasks take text prompts and specific visual conditions (such as segmentation maps, and human pose maps) as multi-modal inputs, then generate new images that comply with the text and image conditions.
Figure 4: (a) Illustration of the construction process for the GRIT-Entity dataset. We used instance segmentation and repainting methods to acquire a large volume of data. (b) Illustration of the cross-verification strategy used in constructing our web images dataset. For a group photo of Person-A and Person-B, we sampled several images from individual photos of Person-A and Person-B and asked MLLM whether they appear in the group photo. A group photo is retained only if the "Yes" ratio for both Person-A and Person-B meets a specific threshold. The individual images marked as "Yes" are then used to construct data pairs with the corresponding group image.
@@ 3.2.2 Subject-driven Image Generation
We constructed both a large-scale foundational dataset (GRIT-Entity dataset) and a high-quality advanced dataset (Web Images dataset) for subject-driven image generation. For the GRIT-Entity dataset, we leveraged the GRIT dataset [51], which annotates object names within images. Using these annotations, we applied the Grounding DINO model [40] for text-to-bounding-box grounding. Based on the bounding boxes, we employed SAM [29] to segment the cropped images, obtaining object masks. We further used the MS-Diffusion model [64] to repaint the object images, enhancing data quality. The process of data construction and the final instruction format are illustrated in Figure˜4-(a). Through this method, we acquired 6 million pairs.
Although the GRIT-based approach provides a substantial amount of data, the input data extracted directly from original images can lead the model to fall into simple copy-paste patterns. To fully unleash the subject-driven image generation capability of OmniGen, we constructed a high-quality web images training dataset using natural images of well-known individuals. First, we sampled 20 million Alt-text entries from the Datacomp dataset [14] and used spaCy [^1] https://github.com/explosion/spaCy for named entity recognition. We selected the most frequently occurring names and employed GPT-4o to filter out real, notable individuals, resulting in 2,000 names. Furthermore, we expanded the initial 2,000 names by including closely related individuals, resulting in approximately 10,000 name pairs. We then scraped images of these individuals and pairs from search engines. Due to the noise in web images, where scraped images may not contain the specified individuals, we designed a cross-verification strategy using InternVL [11] to filter single and group images, as detailed in Figure˜4-(b). The retained single and group images were then captioned with details such as attire and actions. Through additional instruction-based annotations, we successfully constructed a dataset of 533,000 image pairs. We present some examples in Figure 3-(c).

### v2

@@ Abstract
The emergence of Large Language Models (LLMs) has unified language generation tasks and revolutionized human-machine interaction. However, in the realm of image generation, a unified model capable of handling various tasks within a single framework remains largely unexplored. In this work, we introduce OmniGen, a new diffusion model for unified image generation. OmniGen is characterized by the following features: 1) Unification: OmniGen not only demonstrates text-to-image generation capabilities but also inherently supports various downstream tasks, such as image editing, subject-driven generation, and visual-conditional generation. 2) Simplicity: The architecture of OmniGen is highly simplified, eliminating the need for additional plugins. Moreover, compared to existing diffusion models, it is more user-friendly and can complete complex tasks end-to-end through instructions without the need for extra intermediate steps, greatly simplifying the image generation workflow. 3) Knowledge Transfer: Benefit from learning in a unified format, OmniGen effectively transfers knowledge across different tasks, manages unseen tasks and domains, and exhibits novel capabilities. We also explore the model’s reasoning capabilities and potential applications of the chain-of-thought mechanism. This work represents the first attempt at a general-purpose image generation model, and we will release our resources at https://github.com/VectorSpaceLab/OmniGen to foster future advancements.
@@ 1 Introduction
Figure 1: OmniGen can flexibly follow instructions to complete various tasks, without the need for any additional plugins. For complex tasks (e.g., examples in the second line), it can also be completed end-to-end without cumbersome intermediate steps.
The pursuit of Artificial General Intelligence (AGI) has intensified the demand for generative foundation models capable of handling various tasks within a single framework. In the field of Natural Language Processing (NLP), Large Language Models (LLMs) have become exemplary in achieving this goal, demonstrating remarkable versatility across numerous language tasks. However, the realm of visual generation has yet to reveal a counterpart that matches the universality of LLMs. Current image generation models have demonstrated proficiency in specialized tasks. For instance, in the text-to-image generation filed, models such as the SD series [55, 51, 12], DALL-E [54], and Imagen [25] have made significant strides. Meanwhile, lots of efforts have been made to extend the capabilities of diffusion models for specific tasks, such as ControlNet [75], InstandID [64], and InstructPix2Pix [3]. Currently, for each new task, designing a specific module and fine-tuning it is necessary, which hinders the development of image generation. What’s worse, these task-specific models lead to a significant issue: we cannot accomplish tasks simply through input instructions but require a complex and cumbersome workflow involving multiple models and numerous steps. For example, ControlNet needs to use a detector to extract conditions (such as pose estimation maps) and then loads the corresponding condition module. InstantID can only process single-person images and requires a face detection model to detect faces beforehand, followed by using a face encoder to encode the face. If users want to edit the generated images, additional models like InstructPix2Pix need to be loaded.
Can a single model complete various tasks end-to-end through user instructions, similar to how ChatGPT handles language tasks? We envision a future where image generation is made simple and flexible: that any tasks can be accomplished directly through user instructions. Motivated by this goal, we propose a unified framework: OmniGen. As shown in Figure 1, this framework allows for convenient and flexible image generation for any purposes, where no additional plugins or operations are needed. Given the flexibility in following arbitrary instructions, the new framework also help to inspire more interesting image generation tasks.
@@
Unlike popular diffusion models, OmniGen features a very concise structure, comprising only two main components: a VAE and a transformer model. OmniGen supports arbitrarily interleaved text and image inputs as conditions to guide image generation, instead of only accepting pure text or image. To train this architecture as a unified model, we construct a large-scale multi-task image generation dataset X2I, which unifies different tasks with one uniform format. We evaluate the well-trained model based on multiple benchmarks, demonstrating its superior generation capability compared to existing models. Remarkably, OmniGen enables effective knowledge transfer across different scenarios, allowing it to handle unseen tasks and domains while also fostering the emergence of new abilities.
Our contributions are summarized as follows:
@@
  • We introduce OmniGen, a unified image generation model that excels in multiple domains. The model demonstrates competitive text-to-image generation capabilities and inherently supports a variety of downstream tasks, such as controllable image generation and classic computer vision tasks. Furthermore, it can handle complex tasks end-to-end without any lengthy intermediate steps. To our knowledge, OmniGen is the first image generation model to achieve such comprehensive functionality.
@@
  • We construct the first comprehensive image generation dataset named X2I, which stands for ”anything to image”. This dataset covers a wide range of image generation tasks, all of which are standardized into a unified format.
@@
  • By unified training on the multi-task dataset, OmniGen can apply learned knowledge to tackle unseen tasks and domains, as well as exhibit new capabilities. Additionally, we explore the model’s reasoning capabilities and chain-of-thought mechanism.
  • OmniGen takes an initial step toward a foundational model for general image generation. We will open-source the relevant resources (model, code, and data) and hope this work can provide some insights for future image generation models.
@@ 2.1 Model Design
Network Architecture. The design principles of OmniGen are universality and conciseness. As illustrated in Figure 2, OmniGen adopts an architecture comprised of a Variational Autoencoder (VAE) [27] and a pre-trained large transformer model. Specifically, we use the VAE from SDXL [51] to extract continuous visual features from images. The transformer model initialized by Phi-3 [1] generates images based on instructions that serve as conditions. Only VAE is frozen during training. Unlike state-of-the-art diffusion models that require additional encoders to pre-process conditional information (such as CLIP text encoder and image encoder), OmniGen inherently encodes conditional inputs by itself, significantly simplifying the pipeline. Furthermore, OmniGen jointly models text and images within a single model, rather than independently modeling different input conditions with separate encoders as in existing works [68, 70, 72, 64, 8] which lacks interaction between different modality conditions.
@@
Input Representation. The input to the model can be multi-modal interleaved text and images in free form. We utilize the tokenizer of Phi-3 to process text. For images, we firstly employ a VAE with a simple linear layer to extract latent representations. Then, they are flattened into a sequence of visual tokens by linearly embedding each patch in the latent space. Following [49], the patch size is set to 2. We apply standard frequency-based positional embeddings to visual tokens, and process images with varying aspect ratios as SD3 [12]. In addition, we encapsulate each image sequence with two special tokens: “$<$img$>$” and “$<$/img$>$” before inserting it into the sequence of text tokens. We also append the timestep embedding [49] at the end of the input sequence. We do not need any task-specific special tokens to indicate the task type.
@@
Attention Mechanism. Different from text, which can be decomposed into discrete tokens to model, we argue that the image should be modeled as a whole based on its nature. Therefore, we modify the common causal attention mechanism in LLM by integrating it with the bidirectional attention as illustrated in Figure 2. Specifically, we apply causal attention to each element in the sequence, but apply bidirectional attention within each image sequence. This allows each patch to pay attention to other patches within the same image, while ensuring that each image can only attend to other images or text sequences that have appeared previously.
@@
Inference. During inference, we randomly sample a Gaussian noise and then apply the flow matching method to predict the target velocity, iterating multiple steps to obtain the final latent representation. Finally, we use a VAE decoder to decode the latent representation into the predicted image. Thanks to the attention mechanism, OmniGen can accelerate inference like LLMs by using kv-cache: storing previous and current key and value states of the input conditions on the GPU to compute attention without redundant computations.
@@ 2.2 Training Strategy
Train objective. In this work, we use rectified flow [39] to optimize the parameters of model. Different from the DDPM [24], flow matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the step $t$, $\mathbf{x}_{t}$ is defined as: $\mathbf{x}_{t}=t\mathbf{x}+(1-t)\boldsymbol{\epsilon}$, where $\mathbf{x}$ is the original data, and $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised data $\mathbf{x}_{t}$, timestep $t$, and condition information $c$. Specifically, the objective is to minimize the mean squared error loss:
@@
For image editing tasks, the objective is to modify specific regions of the input image while keeping other areas unchanged. Therefore, the difference between the input image and the target image is often small, which allows the model to learn an unexpected shortcut: simply copying the input image as the output to make the related training loss very low. To mitigate this phenomenon, we amplify the loss in the regions of the image where changes occur. More specifically, we calculate the loss weights for each region based on latent representations of input image $\mathbf{x^{\prime}}$ and target image $\mathbf{x}$:
@@
Consequently, regions with alterations are assigned significantly higher weights than those without changes, guiding the model to focus on the areas to be modified.
Training Pipeline. Following previous work [12, 16, 5], we gradually increase the image resolution during the training process. Low resolution is data-efficient, while high resolution can enhance the aesthetic quality of the generated images. Detailed information for each training stage is presented in Table 1. We adopt the AdamW [40] with $\beta=(0.9,0.999)$ as the optimizer. All experiments are conducted on 104 A800 GPUs.
Stage Resolution Steps (K) Batch Size Learning Rate
@@ 5 Multiple 80 104 2e-5
Table 1: Details about each training stage of OmniGen.
@@ 3 X2I Dataset
To achieve robust multi-task processing capabilities, it is essential to train the model on large-scale and diverse datasets. However, in the field of image generation, a readily available large-scale and diverse dataset has yet to emerge. In this work, we have constructed a large-scale unified image generation dataset for the first time, which we refer to as the X2I dataset, meaning ”anything to image”. We have converted all data into a unified format, and Figure 3 presents some examples of the X2I dataset. The entire dataset comprises approximately 0.1 billion images. A detailed description of the composition will be provided in the following sections.
Figure 3: Examples of X2I dataset. We standardized the input of all tasks into an interleaved image-text sequence format. The placeholder ||||image_i|||| represents the position of the $i$-th input image.
@@ 3.1 Text to Image
The input for this subset of data is plain text. We collect multiple open-source datasets from various sources: Recap-DataComp [33](a subset of 56M images), SAM-LLaVA [5], ShareGPT4V [6], LAION-Aesthetic [57](a subset of 4M images), ALLaVA-4V [4], DOCCI [45], DenseFusion [34] and JourneyDB [59]. While these datasets are large in quantity, their image quality is not always high enough. In the early stages of training, we use them to learn a broad range of image-text matching relationships and diverse knowledge. After stage 3, we utilize an internal collection of 16M high-quality images to enhance the aesthetic quality of the generated images. We use the InternVL2 [10] to create synthetic annotations for internal data and LAION-Aesthetic.
@@ 3.2 Multi-modal to Image
Different from existing diffusion models, OmniGen can accept more flexible multi-modal instruction, thus adapting to a variety of data and task types.
@@ 3.2.1 Common Mixed-modal Prompts
The input of this portion of data is arbitrarily interleaved text and images. We collect the data from multiple tasks and sources: image editing (MagicBrush [74], InstructPix2Pix [3] and SEED-edit [17], ), human motion (Something-Something [21]), virtual try-on (HR-VITON [30] and FashionTryon [77]), and style transfer (StyleBooth [23]). The issue of utilizing additional visual conditions for finer-grained spatial control has garnered widespread attention[75, 31]. We employ the MultiGen [52] dataset to learn this function, and select six representative visual conditions: Canny, HED, Depth, Skeleton, Bounding Box, and Segmentation. These types of tasks take text prompts with specific visual conditions (such as segmentation maps, and human pose maps) as multi-modal inputs. We standardize all tasks into the input-output pair format as shown in Figure 3-(b).
@@ 3.2.2 Subject-driven Image Generation
We construct both a large-scale foundational dataset (GRIT-Entity dataset) and a high-quality advanced dataset (Web Images dataset) for subject-driven image generation. For the GRIT-Entity dataset, we leverage the GRIT dataset [50], which annotates object names within images. Using these annotations, we apply GroundingDINO [38] for open-set text-to-bounding-box detection. Based on the grounded bounding boxes, we employ SAM [28] to segment object masks from cropped images. We further use the MS-Diffusion [66] to repaint the object images for higher quality and diversity of data. The process of data construction and the final instruction format are illustrated in supplementary material. Through these methods, we acquire 6M data pairs.
Although the GRIT-based approach provides a substantial amount of samples, the input data extracted directly from original images will tend to lead the model to fall into a simple copy-and-paste pattern. To fully unleash the subject-driven image generation capability of OmniGen, we construct a high-quality natural images dataset from the web. We use the search engine to collect multiple different images of the same person, using one as the input and another as the output. An example is shown in Figure 3-(c). More details about the construction process are given in the supplementary material.